# SSR: Alignment-Aware Modality Connector for Speech Language Models

### v1

@@ Appendix B Evaluation Metric and Prompt
Appendix C Speech MMLU Evaluation
We present the detailed comparison results in Table 9 for better comparison of model performance across different domains / topics. We see that the trend for different domains are mostly consistent, with our alignment-aware connector based on UnitY2 achieving the best performance, followed by Char-CTC based connector. Similar as our main findings, the unit-based system has worse performance due to information loss from discretization and the fine-tuned model suffers from catastrophic forgetting (albeit mitigated through our multitask fine-tuning approach). Nevertheless, all these SSR-Connector based system obtains better performance compared to SpiritLM (LLaMa3), confirming the effectiveness of our modality-fusion strategy.

### v2

@@ Appendix D Evaluation Metric and Prompt
Speech MMLU Evaluation
We craft speech MMLU by synthesizing the questions of MMLU into audio through AudioBox. Since some domains have bad synthesis quality (such as algebra, which includes many math notations), we filtered those domains out from our evaluation.
We present the detailed comparison results in Table 11 for a better comparison of model performance across different domains/topics. We see that the trend for different domains is mostly consistent, with our alignment-aware connector based on UnitY2 achieving the best performance, followed by Char-CTC based connector. Similar as our main findings, the unit-based system has worse performance due to information loss from discretization and the fine-tuned model suffers from catastrophic forgetting (albeit mitigated through our multitask fine-tuning approach). Nevertheless, all these SSR-Connector based system obtains better performance compared to SpiritLM (LLaMa3), confirming the effectiveness of our modality-fusion strategy.